# Predictions for new physics in the CMB damping tail

### v1

@@ Abstract
Ever since the Planck satellite measured the the cosmic microwave background (CMB) down to arcminute angular scales the mismatch between the CMB-inferred value of the Hubble constant and the value inferred from the distance ladder (i.e., the Hubble tension) has been a growing concern and is currently at the $\sim 6\sigma$ level. There are a handful of proposed mechanisms operating in the early universe which have shown some promise in resolving the Hubble tension. These mechanisms are expected to leave a measurable impact on the smallest scale CMB anisotropy, deep in the damping tail. Using current CMB data, baryonic acoustic oscillation data, and the luminosities of type Ia supernovae as a baseline, we compute the predicted small-scale CMB power spectra for a characteristic set of these models. We find that near-future CMB data should be able to detect some but not all of the investigated models.
@@ I Introduction
For around three decades the $\Lambda$CDM model has been firmly established as the baseline cosmological model. It consists of baryons, photons, neutrinos, cold dark matter, and a cosmological constant ($\Lambda$) required to explain the late time accelerated expansion of the universe. During this time it has been repeatedly subjected to a variety of cosmological data and throughout emerged as a consistent description. Various different cosmological probes, such as the cosmic microwave background (CMB) anisotropies, the clustering of large scale structure (LSS) and in particular the baryon acoustic oscillations (BAO), the type Ia supernovae (SNeIa) brightness, cosmic chronometers (CC), as well as the primordial abundance of light elements generated during big bang nucleosynthesis (BBN), have progressively increased the precision of the determination of cosmological parameters to the percent level, see for example Ref. Peebles (2020). However, while these ever tighter measurements continue to be well-fit in the $\Lambda$CDM model, different data sets prefer discrepant values of some of the cosmological parameters. The most significant of these tensions is the difference between the Hubble constant, $H_{0}$ , as measured through late-time direct kinematical measurements and as indirectly inferred through early-time observations. While the value of $H_{0}$ measured through the calibrated SNeIa distance ladder has consistently remained higher than $70$ km/s/Mpc (see e.g. Brout et al. (2022); Riess et al. (2022); Freedman (2021)), recent observations of ever-smaller scales in the CMB anisotropies from the Planck satellite Aghanim et al. (2020a) have shifted the CMB-inferred value downwards to around $68$ km/s/Mpc with a tight uncertainty of $0.5$ km/s/Mpc (the exact value depends slightly on which datasets are used). For about a decade now the possible systematics of either measurement have been scrutinized, without revealing any single systematic that could be responsible for this tension. More worryingly, multiple different probes now cluster in two ‘camps’ of either high or low preferred value of $H_{0}$ Verde et al. (2024). While direct kinematical probes of the recent expansion history (such as from the distance ladder calibrated on Cepheids, Miras, other variable stars, or the TRGB, as well as maser and strong lensing measurements) generally prefer a higher value of the Hubble constant, the indirect probes relying on a dynamical model of the universe (such as the CMB anisotropies, the combination of BAO and BBN, as well as more generally the clustering of the LSS) generally prefer lower values of the Hubble constant. This striking preference of a roughly $\sim 7\%$ larger $H_{0}$ value between different families of probes Verde et al. (2024) clearly motivates the need to look for physics beyond $\Lambda$CDM.
@@
A large number of such models have been proposed, see Schöneberg et al. (2022); Di Valentino et al. (2021); Khalife et al. (2024), but only a few of them have been relatively successful, and none can claim that they unambiguously resolve the tension. Modifications to the post-recombination, ‘late-time’, universe are disfavored due to the fact that once SNeIa are calibrated using any method in the distance ladder, their luminosity distance can be compared to that of the BAO calibrated using the usual sound horizon from the CMB, leading to a clear disagreement over the same range of redshifts Poulin et al. (2024). Indeed, late-time solutions generally struggle with the multitude of observations that prefer an expansion history close to that of the $\Lambda$CDM model Schöneberg et al. (2022). This points towards models which reduce the value of the sound horizon inferred from the CMB, and therefore modify physics at or before recombination.
@@
If we hope to establish that any of these models successfully address the Hubble tension, future observations must be able to clearly establish significant deviations from $\Lambda$CDM. For example, it is well known that additional free-streaming dark radiation causes a phase shift in the acoustic peaks of the CMB and increases the amount of diffusion damping leading to a suppression in power above multipoles of about $\ell\sim 2000$ Aghanim et al. (2020a). However, since any successful model must fit current CMB data, which is itself well-fit by $\Lambda$CDM, we know that a (likely complex) set of parameter degeneracies must come into play which makes identifying simple signatures of any given model difficult. Axion-like EDE is a case in point: ACT DR4 showed a significant preference for this model, along with a significant increase in the CMB-inferred value of $H_{0}$, but it was unclear what exact features which drove this preference Hill et al. (2022); Poulin et al. (2021); Smith et al. (2022a). As we show, the impact of most of these models on multipoles greater than $\ell\sim 2000$ – at scales that are not as well measured by Planck – shows clear and simple deviations from $\Lambda$CDM. The ability to measure these deviations will depend on how well systematic errors from foreground modeling and the non-linear growth of matter perturbations are limited.
We note that there are other approaches to resolving the Hubble tension which may not lead to any strong deviations from $\Lambda$CDM predictions in the CMB. For example, a re-scaling of the gravitational free-fall rates and photon-electron scattering rate leaves most dimensionless cosmological observables nearly invariant, leaving the Hubble constant relatively unconstrained from CMB measurements alone Cyr-Racine et al. (2022). Such a scenario may be realized by proposing a ‘mirror world’ dark sector. The initial scenario presented in Ref. Cyr-Racine et al. (2022) does not work since it transfers the Hubble tension to a tension in the primordial helium abundance but further model building (e.g., Ref. Greene and Cyr-Racine (2024)) may prove to be interesting. Of course these models make other ‘beyond the standard model’ predictions, which will allow them to be tested with future data. A similar degeneracy has been discussed in the context of electron mass variations in Schöneberg and Vacher (2025), which might be probed by large-scale structure data.
@@ II Method
  • It is important to state that this is a conservative estimation, since it might be that a model can still be preferred as only a partial solution to the Hubble tension despite seemingly being ‘excluded’ by future data. In this case the CMB would exclude a full resolution of the Hubble tension but might still be compatible with a partial one.
  • Additionally, there can be shifts in other cosmological parameters (such as $\Omega_{m}$, $n_{s}$, or $\Omega_{b}h^{2}$) that are slightly disfavored by current data but favored by future data. This could either boost or reduce the constraining power for the models we investigate.
  • We purposefully use the somewhat older BAO data (see e.g. Schöneberg et al. (2022)) that was used in the Planck legacy analysis Aghanim et al. (2020a) in order to have a more direct comparison with the corresponding data and best fit files. However, we do expect newer BAO data (such as from the recently released completed eBOSS survey Alam et al. (2021) as well as those from DESI Adame et al. (2025)) to cause small parameter shifts. Given the slight discrepancies between SDSS and DESI in extended parameter space, we decide to use older (and less constraining) BAO data. We use the Pantheon+Brout et al. (2022) SNeIa data, and note that the use of DES-Y5 Abbott et al. (2024) or Union-3 Rubin et al. (2023) would result in small shifts to parameter values due to their preference for different values of $\Omega_{m}$
  • The impact of non-linear structure formation has to be crucially taken into account, as otherwise the analyses would be severely biased. We discuss our approach to this issue in appendix A.
@@
  • Our estimates of the sensitivity of near-future CMB measurements do not account for uncertainties due to modeling of foreground contamination. As such, the error bars present an idealized lower-limit.
@@
We use class as the underlying Einstein-Boltzmann code to predict the power spectra. We note that since class version $>3.2.3$ the in-built precision has been increased by the developers (using a new way of computing the lensing power spectrum based on the Limber approximation that becomes increasingly more accurate at smaller scales), but for older versions the precision needs to be manually increased by increasing the k_max_tau0_over_lmax precision parameter (in our case to $>15$). In either case, despite predicting the lensed CMB power spectra only to $\ell=5000$ it is crucial for accurate lensing predictions to predict the lensing power spectrum at least to $\ell=8000$ (in our case using the parameters l_max_scalars=7000, and the default delta_l_max_lensing=1000). We also set accurate_lensing=1 and l_switch_limber=9. Note that for the MCMC runs we use the low-precision (and hence quicker) default settings of these codes, only requiring high-precision settings when performing the predictions for near-future CMB observations.
For the MCMC runs we use the MontePython code Audren et al. (2013); Brinckmann and Lesgourgues (2018), interfaced with several CMB likelihoods that are already released or are to be released in the near future. Our core analysis uses the plik likelihood analyzing the PR3 data Aghanim et al. (2020b). In order to explore if using PR4 data (NPIPE) Akrami et al. (2020) has an impact, we also analyzed our results using the CamSpec likelihood Rosenberg et al. (2022).
Though, as we discuss in appendix C, we find that this has a mostly negligible impact on the small-scale predictions, and therefore focus on the PR3 predictions here. We use the BAO data from BOSS DR12 Alam et al. (2017), MGS Ross et al. (2015), and 6dFGS Beutler et al. (2011) and the Pantheon+ SNeIa data from Brout et al. (2022). We use the default priors for each likelihood, as well as flat unbounded priors for each of the cosmological parameters, enforcing only $\tau_{\mathrm{reio}}>0.004$. As in Aghanim et al. (2020a) we use a single massive neutrino of $0.06\mathrm{eV}$ mass. The priors on the parameters which are specific to each of the non-$\Lambda$CDM models are largely uninformative.
@@ III Results
We investigate three main models in this work. These are chosen as the most successful models investigated in Ref. Schöneberg et al. (2022) (including variations on their baseline mechanisms), following the three categories outlined in the Introduction. For each of the models we show the predicted power spectra with a prior on the type Ia supernovae absolute magnitude, $M_{b}$ (equivalent to one on $H_{0}$), compared to the reference [^1] The $\Lambda$CDM reference cosmology has $\Omega_{b}h^{2}=0.02242$, $\Omega_{\mathrm{cdm}}h^{2}=0.11933$, $H_{0}=67.66\mathrm{km/s/Mpc}$, $\tau_{\mathrm{reio}}=0.0561$, $A_{s}=2.105\cdot 10^{-9}$ and $n_{s}=0.9665$, which are the officially reported mean values in Aghanim et al. (2020a). $\Lambda$CDM power spectrum obtained without a prior on $M_{b}$ Explicitly, we define the difference $\Delta C_{\ell}^{\rm XY}\equiv(C_{\ell}^{\rm XY})_{{\rm new\ model},M_{b}}-(C_{\ell}^{\rm XY})_{\Lambda{\rm CDM},\mathrm{no}\leavevmode\nobreak\ M_{b}}$ for the supplied figures. We randomly draw 100 points from the part of the MCMC chain that has the burn-in removed in order to draw the mean and standard deviation lines/contours; we note that increasing this number to 1000 points did not significantly change the results.
@@ III.1 Wess-Zumino dark radiation
As discussed before, in order to increase the CMB-inferred Hubble constant it is necessary to decrease the sound horizon Schöneberg et al. (2022); Poulin et al. (2024). The introduction of ‘dark radiation’ – radiation which effectively does not interact with visible matter – is a natural choice. There are many possible extensions of the core model of particle physics which can achieve this, chief among them an increase in the total energy density in neutrinos. However, such a change is significantly limited by the fact that neutrinos free-stream, leading to a measurable phase shift in the acoustic oscillations Bashinsky and Seljak (2004); Follin et al. (2015); Baumann et al. (2019); Saravanan et al. (2025). On the other hand, strongly self-interacting dark radiation models do not lead to changes in the phase as well as partially compensating the increased diffusion damping due to enhanced clustering on smaller scales (see (Blinov and Marques-Tavares, 2020, Sec. 3.2)) and are currently a viable way to address the Hubble tension. However, as is the case with all models that introduce additional energy density in the pre-recombination universe, these models significantly impact the diffusion damping scale Brust et al. (2017); Blinov and Marques-Tavares (2020); Becker et al. (2021); Archidiacono et al. (2019); Saravanan et al. (2025).
The downside of these models is that the amount of dark radiation required to significantly reduce the Hubble tension is relatively large (around $\Delta N_{\mathrm{eff}}\simeq 0.7-0.8$). Such a large amount of extra dark radiation significantly impacts the diffusion damping scale almost independently of the nature of the dark radiation Brust et al. (2017); Blinov and Marques-Tavares (2020); Becker et al. (2021); Archidiacono et al. (2019); Saravanan et al. (2025).
In addition to considering strongly self-interacting dark radiation, one can also change when its abundance is generated. Often times this is done either through decay of some heaver particle or through thermalization of the dark radiation bath with another heat bath. This mechanism of changing the abundance is often times employed when considering why dark radiation would not lead to significant BBN constraints, as an increase in the Hubble parameter at early times around BBN would lead to less time for the neutrons to decay before forming Helium and therefore would lead to a higher Helium abundance that is not observed. However, the same mechanism can also be employed for the CMB itself. If the time during which the dark radiation abundance is increased is set at a redshift of around $z\gtrsim 10^{4}$ (about a decade in redshift before recombination), modes on the smaller scales ($\ell\gtrsim 1000$) that are very sensitive to diffusion damping will have experienced a smaller amount of dark radiation when entering the Hubble horizon, while modes on the larger scales ($\ell\lesssim 1000$) which are also those scales for which the sound horizon can be very well measured will have experienced a larger amount of dark radiation. Overall, this leads to the possibility of having a small sound horizon with a comparatively smaller amount of diffusion damping experienced on the smallest scales Schöneberg and Franco Abellán (2022).

### v2

@@ Abstract
Ever since the Planck satellite measured the the cosmic microwave background (CMB) down to arcminute angular scales, the mismatch between the CMB-inferred value of the Hubble constant and the value inferred from the distance ladder (i.e., the Hubble tension) has been a growing concern and is currently at the $\sim 6\sigma$ level. There are a handful of proposed mechanisms operating in the early universe which have shown some promise in resolving the Hubble tension. These mechanisms are expected to leave a measurable impact on the smallest scale CMB anisotropy, deep in the damping tail. Using current CMB data, baryonic acoustic oscillation data, and the luminosities of Type Ia supernovae as a baseline, we compute the predicted small-scale CMB power spectra for a characteristic set of these models. We find that near-future CMB data should be able to distinguish some but not all of the investigated models from the core cosmological model, $\Lambda$CDM.
@@ I Introduction
For around three decades the $\Lambda$CDM model has been firmly established as the baseline cosmological model. It consists of baryons, photons, neutrinos, cold dark matter, and a cosmological constant ($\Lambda$) required to explain the late time accelerated expansion of the universe. During this time it has been repeatedly subjected to a variety of cosmological data and throughout emerged as a consistent description. Various different cosmological probes, such as the cosmic microwave background (CMB) anisotropies, the clustering of large scale structure (LSS) and in particular the baryon acoustic oscillations (BAO), the Type Ia supernovae (SNeIa) brightness, cosmic chronometers (CC), as well as the primordial abundance of light elements generated during big bang nucleosynthesis (BBN), have progressively increased the precision of the determination of cosmological parameters to the percent level, see for example Ref. Peebles (2020). However, while these ever tighter measurements continue to be well-fit in the $\Lambda$CDM model, different data sets prefer discrepant values of some of the cosmological parameters. The most significant of these tensions is the difference between the Hubble constant, $H_{0}$ , as measured through late-time direct kinematical measurements and as indirectly inferred through early-time observations. While the value of $H_{0}$ measured through the calibrated SNeIa distance ladder has consistently remained higher than $70$ km/s/Mpc (see e.g. Ref. Brout et al. (2022); Riess et al. (2022); Freedman (2021)), recent observations of ever-smaller scales in the CMB anisotropies from the Planck satellite Aghanim et al. (2020a) have shifted the CMB-inferred value downwards to around $68$ km/s/Mpc with a tight uncertainty of $0.5$ km/s/Mpc (the exact value depends slightly on which datasets are used).
The possible systematics of either measurement have been scrutinized, without revealing any single systematic that could be responsible for this tension (see, e.g., Refs. Riess et al. (2022); Abdalla et al. (2022)). More worryingly, multiple different probes now cluster in two ‘camps’ of either high or low preferred value of $H_{0}$ Verde et al. (2024). While direct kinematical probes of the recent expansion history (such as from the distance ladder calibrated on Cepheids, Miras, other variable stars, or the TRGB, as well as maser and strong lensing measurements) generally prefer a higher value of the Hubble constant, the indirect probes relying on a dynamical model of the universe (such as the CMB anisotropies, the combination of BAO and BBN, as well as more generally the clustering of the LSS) generally prefer lower values of the Hubble constant. This striking preference of a roughly $\sim 7\%$ larger $H_{0}$ value between different families of probes Verde et al. (2024) clearly motivates the need to look for physics beyond $\Lambda$CDM.
@@
A large number of such models have been proposed, see Schöneberg et al. (2022); Di Valentino et al. (2021); Khalife et al. (2024), but only a few of them have been relatively successful, and none can claim that they unambiguously resolve the tension. Modifications to the post-recombination, ‘late-time’, universe are disfavored due to the fact that once SNeIa are calibrated using any method in the distance ladder, their luminosity distance can be compared to that of the BAO calibrated using the usual sound horizon from the CMB, leading to a clear disagreement over the same range of redshifts Efstathiou (2021); Raveri (2023); Poulin et al. (2024). Indeed, late-time solutions generally struggle with the multitude of observations that prefer an expansion history close to that of the $\Lambda$CDM model Schöneberg et al. (2022). This points towards models which reduce the value of the sound horizon inferred from the CMB, and therefore modify physics at or before recombination.
@@
If any of these models successfully address the Hubble tension, future observations must be able to clearly establish significant deviations from $\Lambda$CDM. For example, it is well-known that additional free-streaming dark radiation causes a phase shift in the acoustic peaks of the CMB and increases the amount of diffusion damping leading to a suppression in power above multipoles of about $\ell\sim 2000$ Aghanim et al. (2020a). However, since any successful model must fit current CMB data, which is itself well-fit by $\Lambda$CDM, we know that a (likely complex) set of parameter degeneracies must come into play which makes identifying simple signatures of any given model difficult. Axion-like EDE is a case in point: ACT DR4 showed a significant preference for this model, along with a significant increase in the CMB-inferred value of $H_{0}$, but it was unclear what exact features which drove this preference Hill et al. (2022); Poulin et al. (2021); Smith et al. (2022a). As we show, the impact of most of these models on multipoles greater than $\ell\sim 2000$ – at scales that are not as well measured by Planck – leads to clear and simple deviations from $\Lambda$CDM. The ability to measure these deviations will depend on how well systematic errors from foreground modeling and the non-linear growth of matter perturbations are limited.
We note that there are other approaches to resolving the Hubble tension which may not lead to any strong deviations from $\Lambda$CDM predictions in the CMB. For example, a re-scaling of the gravitational free-fall rates and photon-electron scattering rate leaves most dimensionless cosmological observables nearly invariant, leaving the Hubble constant relatively unconstrained from CMB measurements alone Cyr-Racine et al. (2022); Baryakhtar et al. (2024). Such a scenario may be realized by proposing a ‘mirror world’ dark sector. The initial scenario presented in Ref. Cyr-Racine et al. (2022) does not work since it transfers the Hubble tension to a tension in the primordial helium abundance but further model building (e.g., Ref. Greene and Cyr-Racine (2024)) may prove to be interesting. Of course these models make other ‘beyond the standard model’ predictions, which will allow them to be tested with future data. A related degeneracy has been discussed in the context of electron mass variations in Schöneberg and Vacher (2025), which can be probed by large-scale structure data.
@@ II Method
  • It is important to state that this is a conservative estimate, since it might be that a model can still be preferred as only a partial solution to the Hubble tension despite seemingly being ‘excluded’ by future data. In this case the CMB would exclude a full resolution of the Hubble tension but might still be compatible with a partial one.
  • Additionally, there may be shifts in other cosmological parameters (such as $\Omega_{m}$, $n_{s}$, or $\Omega_{b}h^{2}$) that are slightly disfavored by current data but favored by future data. This could either boost or reduce the constraining power for the models we investigate.
  • We purposefully use the somewhat older BAO data (see e.g. Ref. Schöneberg et al. (2022)) that was used in the Planck legacy analysis Aghanim et al. (2020a) in order to have a more direct comparison with the corresponding data and best fit files. However, we do expect newer BAO data (such as from the recently released completed eBOSS survey Alam et al. (2021) as well as those from DESI Abdul Karim et al. (2025)) to cause small parameter shifts. Given the slight discrepancies between SDSS and DESI in extended parameter space (see, e.g., Ref. Chaussidon et al. (2025)), we decide to use older (and less constraining) BAO data. We use the Pantheon+Brout et al. (2022) SNeIa data, and note that the use of DES-Y5 Abbott et al. (2024) or Union-3 Rubin et al. (2023) would also result in small shifts to parameter values due to their preference for different values of $\Omega_{m}$
  • The impact of non-linear structure formation has to be crucially taken into account, as otherwise the analyses would be severely biased. We discuss our approach to this issue in appendix A (see also Ref. Trendafilova et al. (2025)).
@@
  • Our estimates of the sensitivity of near-future CMB measurements do not account for uncertainties due to modeling of foreground contamination. As such, the error bars present an idealized lower-limit. We also note that foreground modeling can introduce strong correlations between multipoles.
@@
We use class as the underlying Einstein-Boltzmann code to predict the power spectra. We note that since class version $>3.2.3$ the in-built precision has been increased by the developers (using a new way of computing the lensing power spectrum based on the Limber approximation that becomes increasingly more accurate at smaller scales), but for older versions the precision needs to be manually increased by increasing the k_max_tau0_over_lmax precision parameter (in our case to $15$). In either case, despite predicting the lensed CMB power spectra only to $\ell=5000$ it is crucial for accurate lensing predictions to predict the lensing power spectrum at least to $\ell=8000$ (in our case using the parameters l_max_scalars=7000, and the default delta_l_max_lensing=1000). We also set accurate_lensing=1 and l_switch_limber=9. Note that for the MCMC runs we use the low-precision (and hence quicker) default settings of these codes, only requiring high-precision settings when performing the predictions for near-future CMB observations.
For the MCMC runs we use the MontePython code Audren et al. (2013); Brinckmann and Lesgourgues (2018), interfaced with several CMB likelihoods that are already released or are to be released in the near future. Our core analysis uses the plik likelihood analyzing the PR3 data Aghanim et al. (2020b). In order to explore if using Planck PR4 data (NPIPE) Akrami et al. (2020) has an impact, we also analyzed our results using the CamSpec likelihood Rosenberg et al. (2022). Though, as we discuss in appendix C, we find that this has a mostly negligible impact on the small-scale predictions, and therefore focus on the PR3 predictions here. We use the BAO data from BOSS DR12 Alam et al. (2017), MGS Ross et al. (2015), and 6dFGS Beutler et al. (2011) and the Pantheon+ SNeIa data from Brout et al. (2022). We use the default priors for each likelihood, as well as flat unbounded priors for each of the cosmological parameters, enforcing only $\tau_{\mathrm{reio}}>0.004$. As in Aghanim et al. (2020a) we use a single massive neutrino of $0.06\mathrm{eV}$ mass. The priors on the parameters which are specific to each of the non-$\Lambda$CDM models are uninformative.
@@ III Results
We investigate three main models in this work. These are chosen as the most successful models investigated in Ref. Schöneberg et al. (2022) (including variations on their baseline mechanisms), following the three categories outlined in the Introduction. For each of the models we show the predicted power spectra with a prior on the Type Ia supernovae absolute magnitude, $M_{b}$ (equivalent to one on $H_{0}$), compared to the reference [^1] The $\Lambda$CDM reference cosmology has $\Omega_{b}h^{2}=0.02242$, $\Omega_{\mathrm{cdm}}h^{2}=0.11933$, $H_{0}=67.66\mathrm{km/s/Mpc}$, $\tau_{\mathrm{reio}}=0.0561$, $A_{s}=2.105\cdot 10^{-9}$, $n_{s}=0.9665$, and a single massive neutrino of $0.06\mathrm{eV}$ mass, which are the officially reported mean values in Aghanim et al. (2020a). $\Lambda$CDM power spectrum obtained without a prior on $M_{b}$ Explicitly, we define the difference $\Delta C_{\ell}^{\rm XY}\equiv(C_{\ell}^{\rm XY})_{{\rm new\ model},M_{b}}-(C_{\ell}^{\rm XY})_{\Lambda{\rm CDM},\mathrm{no}\leavevmode\nobreak\ M_{b}}$ for the supplied figures. We randomly draw 100 points from the part of the MCMC chain that has the burn-in removed in order to draw the mean and standard deviation lines/contours; we note that increasing this number to 1000 points did not significantly change the results.
@@ III.1 Wess-Zumino dark radiation
As discussed before, in order to increase the CMB-inferred Hubble constant it is necessary to decrease the sound horizon Schöneberg et al. (2022); Poulin et al. (2024). The introduction of ‘dark radiation’ – radiation which effectively does not interact with visible matter – is a natural choice. There are many possible extensions of the core model of particle physics which can achieve this, chief among them an increase in the total energy density in neutrinos. However, such a change is significantly limited by the fact that neutrinos free-stream, leading to a measurable phase shift in the acoustic oscillations Bashinsky and Seljak (2004); Follin et al. (2015); Baumann et al. (2019); Saravanan et al. (2025); Montefalcone et al. (2025). On the other hand, strongly self-interacting dark radiation models do not lead to changes in the phase as well as partially compensate the increased diffusion damping due to enhanced clustering on smaller scales (see (Blinov and Marques-Tavares, 2020, Sec. 3.2)) and are currently a viable way to address the Hubble tension. However, as is the case with all models that introduce additional energy density in the pre-recombination universe, these models significantly impact the diffusion damping scale Brust et al. (2017); Blinov and Marques-Tavares (2020); Becker et al. (2021); Archidiacono et al. (2019); Saravanan et al. (2025).
In addition to considering strongly self-interacting dark radiation, one can also change when its abundance is generated. Often times this is done either through decay of some heaver particle or through thermalization of the dark radiation bath with another heat bath. For example, in order to avoid constraints from measurements of the light element abundances and big bang nucleosynthesis (BBN), the dark radiation may be produced after BBN (e.g., Ref. Aloni et al. (2023)). The same mechanism can also be employed for the CMB itself. If the time during which the dark radiation abundance is increased is set at a redshift of around $z\gtrsim 10^{4}$ (about a decade in redshift before recombination), modes on the smaller scales ($\ell\gtrsim 1000$) that are very sensitive to diffusion damping will have experienced a smaller amount of dark radiation when entering the Hubble horizon, while modes on the larger scales ($\ell\lesssim 1000$) which are also those scales for which the sound horizon can be very well measured will have experienced a larger amount of dark radiation. Overall, this leads to the possibility of having a small sound horizon with a comparatively smaller amount of diffusion damping experienced on the smallest scales Schöneberg and Franco Abellán (2022).